OpenStudy (anonymous):

A sphere has a radius of 2x+5 Which polynomial in standard form best describes the total surface area of the sphere? Use the formula S= 4 pi r^2 for the surface area of a sphere

OpenStudy (anonymous):

4pi(2x+5)^2 (2x+5)(2x+5)

OpenStudy (anonymous):

is this the right answer?

OpenStudy (anonymous):

@KenLJW

OpenStudy (kenljw):

The volume of a sphere is V = (4/3) pi r^3 The surface area of a sphere is A = 4 pi r^2 In your case A = 4 pi (2 x + 5) ^ 2 A = 4 pi (4 x^ 2 + 20 x +25) A = 16 pi (x^ 2 + 5 x + 25/4) is this what you mean by standard form?